OpenStudy (anonymous):

if u= log {x^2+y^2/x+y} prove that x du/dx + y du/dy=1

myininaya (myininaya):

let me try this again lol

myininaya (myininaya):

i'm gonna try something easier

OpenStudy (anonymous):

Okay \[\large{\frac{du}{dx} \frac{1}{ \frac{y^2+x^2 }{x+y}}\frac{(2x + 2y\frac{dy}{dx})(x+y)+(1+\frac{dy}{dx}(x^2+y^2)}{(x+y)^2} }\]

OpenStudy (anonymous):

why are you asking these dipti .....these are not tough ones ..these are ugly ones .

myininaya (myininaya):

this is it i read it wrong at first i think he means partial derivative i used this same notation i should had used correct notation but i used his

OpenStudy (anonymous):

ok i did a typo there in numerator f(x)g'(x) - g(x)f'(x) not +

myininaya (myininaya):

did you mean partial derivative?